OpenStudy (anonymous):

i really need help.1.Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem

OpenStudy (anonymous):

We have to write a system of equations using 3 variables. First things first we need to calculate the sale price which is the purchase price(0.50) plus the purchase price(0.50) times (40% or 0.40) you should get around (0.70 or $0.70). Now for the equations; if we take the $0.50 purchase price times the variable X(number of flowers purchased) you get a variable Y1 (which we will call total price of flowers bought). If we take the the $0.70 sale price times the variable X(number of flowers purchased again) we get Y2(which we will call total price of flowers sold). Now to consider the profit Angelica's service club made, if we take Y2 - Y1 = $500(profit); this is saying essentially that the total price of flowers bought minus the total price of flowers sold should equal the profit. I will write it in equation form for you \[Y1 = X*(0.50) \] \[Y2 = X*(0.70)\] \[Y2 - Y1 = 500\] 3 equations and 3 variable you should be able to solve these simple equations. Then answer should be X = 2500 which is the number of flowers that should be bought.